OpenStudy (anonymous):

Running out of time What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = x +

OpenStudy (anonymous):

y=mx+b so you need to find b so this is that you do (3,4)- 4 goes in for y and 3 goes in for x like this 4=3+b then subtract 3 from 4 and you get 1 and the slope is 2 so y=2x+1

OpenStudy (zpupster):

y=mx+b 4=2(3)+b 4= 6+b 4-6=b -2=b y=2x-2

OpenStudy (anonymous):

Dont you mean + 2?

OpenStudy (danjs):

Point (3,4) slope = 2 point-slope form y - y1 = m(x-x1) y - 4 = 2*(x-3)

OpenStudy (danjs):

put that into form they want, y=mx + b, by distributing the 2, and isolating Y by itself

OpenStudy (anonymous):

y=2x+1 is what i had

OpenStudy (anonymous):

so im right?

OpenStudy (danjs):

y - 4 = 2*(x-3) y - 4 = 2x - 6 y = 2x - 6 + 4 y = 2x - 2

OpenStudy (anonymous):

uh it wont let me put a - tho, it has a + and i cant change it

OpenStudy (danjs):

+ (-2)

OpenStudy (danjs):

plus a negative, is the same as, minus

OpenStudy (danjs):

y = 2x + (-2)

OpenStudy (anonymous):

Ok i guess... im pretty unsure :/

OpenStudy (danjs):

that is what it is, type -2 for the b spot in y=mx+b

OpenStudy (danjs):

m=2 b = -2